# Simulation-Based Inference for Probabilistic Galaxy Detection and Deblending

### v1

@@ Acknowledgments
We are grateful to the developers of the Python modules that we used: Astropy (astropy:2013; astropy:2018; astropy2022), Blending ToolKit (mendoza2025blending), einops (rogozhnikov2022einops), GalSim (galsim2015), Matplotlib (matplotlib2007), NumPy (numpy2020), Pytest (pytest), PyTorch (pytorch2019), PyTorch Lightning (pytorch_lightning2019), Ruff [^23] https://github.com/astral-sh/ruff, scikit-learn (scikit-learn), SciPy (scipy2020), SEP (bertin1996sextractor; barbary2016sep) [^24] https://github.com/sep-developers/sep, SurveyCodex [^25] https://github.com/LSSTDESC/surveycodex, Tensorboard (tensorflow2015-whitepaper), Typer [^26] https://typer.tiangolo.com/, and tqdm (typer).
@@
The code to reproduce all experiments is available in a public GitHub repository: https://github.com/prob-ml/bliss/tree/desc-oja. The software is also available in Zenodo at https://zenodo.org/records/18164685.
@@ Appendix A Appendix A: Bayesian framework for probabilistic cataloging
Traditionally, Bayesian inference for astronomical catalogs might use the conditional likelihood paired with Markov chain Monte Carlo (MCMC) sampling (schneider2015hierarchical; portillo2017improved; feder2020multiband). Given the light source counts in an image and a parametric source model, one can run an MCMC chain over all sources and source parameters targeting the conditional likelihood $P(\mathcal{X}|\mathcal{Z})$ to output a joint posterior distribution of the sources’ parameters. Although this fully Bayesian approach promises exact posterior samples upon convergence, MCMC sampling is less computationally tractable over larger parameter spaces and data volumes compared with a variational inference approach (blei2017variational).
@@ Appendix B Appendix B: Variational loss function derivation
where $P(\mathcal{Z}|\mathcal{X})$ is the true posterior, $Q_{\bm{\phi}}$ for $Q_{\bm{\phi}}(\mathcal{Z})$ is the variational distribution, $\mathcal{Z}$ are the light source parameters (Equation 7), $\mathcal{X}$ is an astronomical image, and $D_{\rm KL}(\cdot\,\|\,\cdot)$ is the KL divergence.
@@
where $P$ and $Q_{\bm{\phi}}$ are shorthand for $P(\mathcal{Z},\mathcal{X})$ and $Q_{\bm{\phi}}(\mathcal{Z})$, respectively. We can now expand the KL divergence using the definition to obtain:
@@
$=\operatorname*{argmax}_{\bm{\phi}}\mathbb{E}_{\mathcal{X}\sim P(\mathcal{X})}\mathbb{E}_{\mathcal{Z}\sim P(\mathcal{Z}|\mathcal{X})}\log Q_{\bm{\phi}}(\mathcal{Z}).$ (26)
@@
$\bm{\phi}^{\star}$ $=\operatorname*{argmax}_{\bm{\phi}}\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\log Q_{\bm{\phi}}(\mathcal{Z})$ (27)
$=\operatorname*{argmax}_{\bm{\phi}}\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\sum_{t=1}^{T}\log Q_{\bm{\phi}}(\mathcal{Z}_{t}).$ (28)
@@
$L^{\prime}(\bm{\phi})$ $=\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\sum_{t=1}^{T}\log Q_{\bm{\phi}}(n_{t})+\log Q_{\bm{\phi}}(\bm{\ell}_{t}|n_{t})$
$+\log Q_{\bm{\phi}}(f_{t}|n_{t},\bm{\ell}_{t},b_{t})+\log Q_{\bm{\phi}}(b_{t}|n_{t},\bm{\ell}_{t}),$ (30)
@@ Appendix C Appendix C: Settings for SourceExtractor in Python
We use SExtractor in Python (SEP) v1.4.11.4.1 (barbary2016sep). The convolution kernel (FILTER) corresponds to a Gaussian profile in a $3\times 3$ array of pixel values with a width of $2.0$ pixels ($0.4$ arcsec).
The specific arguments used for detection through the sep.extract function are listed in Table C. We aim to match the SExtractor (bertin1996sextractor) settings used in sanchez2021effects as much as possible.

### v2

@@ Acknowledgments
AG acknowledge the support of a grant from the Simons Foundation (Simons Investigator in Astrophysics, Award ID 620789).
We are grateful to the developers of the Python modules that we used: Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Blending ToolKit (Mendoza et al., 2025), einops (Rogozhnikov, 2022), GalSim (Rowe et al., 2015), Matplotlib (Hunter, 2007), NumPy (Harris et al., 2020), Pytest (Krekel et al., 2004), PyTorch (Paszke et al., 2019), PyTorch Lightning (Falcon and others, 2019), Ruff [^24] https://github.com/astral-sh/ruff, scikit-learn (Pedregosa et al., 2011), SciPy (Virtanen et al., 2020), SEP (Bertin and Arnouts, 1996; Barbary, 2016) [^25] https://github.com/sep-developers/sep, SurveyCodex [^26] https://github.com/LSSTDESC/surveycodex, Tensorboard (Abadi et al., 2015), Typer [^27] https://typer.tiangolo.com/, and tqdm (da Costa-Luis et al., 2024).
@@
The code to reproduce all experiments is in the public GitHub repository: https://github.com/prob-ml/bliss/tree/desc-oja. The software is also available in Zenodo at https://zenodo.org/records/18164685.
@@ Appendix A Appendix A: Bayesian framework for probabilistic cataloging
Traditionally, Bayesian inference for astronomical catalogs might use the conditional likelihood paired with Markov chain Monte Carlo (MCMC) sampling (Schneider et al., 2015; Portillo et al., 2017; Feder et al., 2020). Given the light source counts in an image and a parametric source model, one can run an MCMC chain over all sources and source parameters targeting the conditional likelihood $P(\mathcal{X}|\mathcal{Z})$ to output a joint posterior distribution of the sources’ parameters. Although this fully Bayesian approach promises exact posterior samples upon convergence, MCMC sampling is less computationally tractable over larger parameter spaces and data volumes compared with a variational inference approach (Blei et al., 2017).
@@ Appendix B Appendix B: Variational loss function derivation
where $P(\mathcal{Z}|\mathcal{X})$ is the true posterior, $Q_{\bm{\phi}}(\mathcal{Z})$ is the variational distribution, $\mathcal{Z}$ are the light source parameters (Equation 7), $\mathcal{X}$ is an astronomical image, and $D_{\rm KL}(\cdot\,\|\,\cdot)$ is the KL divergence.
@@
where $P$ and $Q_{\bm{\phi}}$ are shorthand for $P(\mathcal{Z}|\mathcal{X})$ and $Q_{\bm{\phi}}(\mathcal{Z})$, respectively. We can now expand the KL divergence using the definition (Equation 9) to obtain:
@@
$=\operatorname*{argmin}_{\bm{\phi}}-\mathbb{E}_{\mathcal{X}\sim P(\mathcal{X})}\mathbb{E}_{\mathcal{Z}\sim P(\mathcal{Z}|\mathcal{X})}\log Q_{\bm{\phi}}(\mathcal{Z}).$ (26)
@@
$\bm{\phi}^{\star}$ $=\operatorname*{argmin}_{\bm{\phi}}-\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\log Q_{\bm{\phi}}(\mathcal{Z})$ (27)
$=\operatorname*{argmin}_{\bm{\phi}}-\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\sum_{t=1}^{T}\log Q_{\bm{\phi}}(\mathcal{Z}_{t}).$ (28)
@@
$L^{\prime}(\bm{\phi})$ $=-\mathbb{E}_{\mathcal{X,Z}\sim P(\mathcal{X,Z})}\sum_{t=1}^{T}\Big[\log Q_{\bm{\phi}}(n_{t})+\log Q_{\bm{\phi}}(\bm{\ell}_{t}|n_{t})$
$+\log Q_{\bm{\phi}}(b_{t}|n_{t},\bm{\ell}_{t})\Big],$ (30)
@@ Appendix C Appendix C: Settings for SourceExtractor in Python
We use SExtractor in Python (SEP) v1.4.11.4.1 (Barbary, 2016). The convolution kernel (FILTER) corresponds to a Gaussian profile in a $3\times 3$ array of pixel values with a width of $2.0$ pixels ($0.4$ arcsec).
The specific arguments used for detection through the sep.extract function are listed in Table C. We aim to match the SExtractor (Bertin and Arnouts, 1996) settings used in Sanchez et al. (2021a) as much as possible.